# Varying coefficients correlated velocity models in complex landscapes with boundaries applied to narwhal responses to noise exposure

### v3

@@ 4.2 Linear Gaussian state-space model
In the baseline statistical model (10), we make the approximation suggested in (Michelot et al., 2021) that the function $\omega_{i}$ is constant on each time step $[t_{ij},t_{i,j+1}]$, equal to its value at time $t_{ij}$, that is
@@
A linear Gaussian state-space model is obtained similarly for the response model. Estimation with smoothSDE relies on this state-space formulation. The complete likelihood is computed from the observations $y_{ij}$ as a by-product of the Kalman filter algorithm (Michelot et al., 2021). Laplace’s approximation of the integral of this complete likelihood over the random effects is computed using the R package TMB and optimization is performed via the optim function in R with the BFGS gradient method, with the gradient being calculated by automatic differenciation. We refer to (Kristensen et al., 2016) for more details about the derivation of the multidimensional Laplace’s approximation and the gradient computations. We refer to (Wood, 2017, section 6.6) for the specification of the mixed effect model and the regularization terms that are based on a wiggliness measure involving Gaussian priors for the spline coefficients $\omega_{k}$, $k\in\{1,\cdots,L\}$.
@@ 6.1 Baseline estimation
We estimate the baseline model as described in Section 4.1 on the data before exposure. We set the degree of freedom of the marginal splines to $5$ leading to a total number of $25$ degrees of freedom, fix the smoothing penalties of the tensor splines to $1$, and initialize the random effects standard deviations to $\sigma_{\nu}=\sigma_{\tau}=1$, and the spline coefficients $\omega_{k}$, $k\in\{1,\cdots,L\}$ to 0. The measurement error standard deviation is fixed to $\sigma_{obs}=50$ m based on the measurement errors of Fastloc-GPS found in (Wensveen, Thomas and Miller, 2015).
Estimation and confidence intervals on the parameter scale are shown in Table 2. The persistence is estimated to $\hat{\tau}_{0}=1.35$ h. In comparison, harbour seal in Alaska were shown to exhibit slightly more persistent motion $\hat{\tau}=1.51$ h with $95\%$ CI ; $1.30-1.75$ (Johnson et al., 2008) while bowhead whales in Greenland showed much less persistence $\hat{\tau}=0.17$ h with $95\%$ CI ; $0.14-0.20$ (Gurarie et al., 2017). The intercept value for $\nu$ is estimated to $\hat{\nu}_{0}=4.60$ km/h which is of the order of magnitude of the average of the observed velocities.
@@ 6.2 Response estimation
We draw special attention to the parameter $\tau$, which decreases with increased exposure to the ship, implying lower persistence and lower autocorrelation in the velocity of the narwhals. This had been hinted at in (Heide-Jørgensen et al., 2021), where it was shown that the narwhals have more tendency to change direction and move towards the shore in presence of the ship. This can be interpreted as a drop in persistence due to the appearance of the ship. Our analysis suggests that the magnitude of this drop might be alarming: for most of the narwhals, at distance less than $5$ km from the ship, the persistence is half the baseline value. This is evidence of a strong shift in the behavior of the whales, which might have consequences on their capacity to rest and forage at short term.
@@ 7 Conclusion and perspectives
The authors would like to thank the CNRS and IRN Madef for funding A. Delporte, and MIAI - ANR-19-P3IA-0003 for the funding of A Samson.

### v5

@@ 4.2 Linear Gaussian state-space model
In the baseline statistical model (10), we make the approximation suggested in (michelot_varying-coefficient_2021) that the function $\omega_{i}$ is constant on each time step $[t_{ij},t_{i,j+1}]$, equal to its value at time $t_{ij}$, that is
@@
A linear Gaussian state-space model is obtained similarly for the response model. Estimation with smoothSDE relies on this state-space formulation. The complete likelihood is computed from the observations $y_{ij}$ as a by-product of the Kalman filter algorithm (michelot_varying-coefficient_2021). Laplace’s approximation of the integral of this complete likelihood over the random effects is computed using the R package TMB and optimization is performed via the optim function in R with the BFGS gradient method, with the gradient being calculated by automatic differenciation. We refer to (kristensen_tmb_2016) for more details about the derivation of the multidimensional Laplace’s approximation and the gradient computations. We refer to (wood_generalized_2017, section 6.6) for the specification of the mixed effect model and the regularization terms that are based on a wiggliness measure involving Gaussian priors for the spline coefficients $\omega_{k}$, $k\in\{1,\cdots,L\}$.
@@ 6.1 Baseline estimation
We estimate the baseline model as described in Section 4.1 on the data before exposure. We set the degree of freedom of the marginal splines to $5$ leading to a total number of $25$ degrees of freedom, fix the smoothing penalties of the tensor splines to $1$, and initialize the random effects standard deviations to $\sigma_{\nu}=\sigma_{\tau}=1$, and the spline coefficients $\omega_{k}$, $k\in\{1,\cdots,L\}$ to 0. The measurement error standard deviation is fixed to $\sigma_{obs}=50$ m based on the measurement errors of Fastloc-GPS found in (wensveen_path_2015).
Estimation and confidence intervals on the parameter scale are shown in Table 2. The persistence is estimated to $\hat{\tau}_{0}=1.35$ h. In comparison, harbour seal in Alaska were shown to exhibit slightly more persistent motion $\hat{\tau}=1.51$ h with $95\%$ CI ; $1.30-1.75$ (johnson_continuous_2008) while bowhead whales in Greenland showed much less persistence $\hat{\tau}=0.17$ h with $95\%$ CI ; $0.14-0.20$ (gurarie_correlated_2017). The intercept value for $\nu$ is estimated to $\hat{\nu}_{0}=4.60$ km/h which is of the order of magnitude of the average of the observed velocities.
@@ 6.2 Response estimation
We draw special attention to the parameter $\tau$, which decreases with increased exposure to the ship, implying lower persistence and lower autocorrelation in the velocity of the narwhals. This had been hinted at in (heide-jorgensen_behavioral_2021), where it was shown that the narwhals have more tendency to change direction and move towards the shore in presence of the ship. This can be interpreted as a drop in persistence due to the appearance of the ship. Our analysis suggests that the magnitude of this drop might be alarming: for most of the narwhals, at distance less than $5$ km from the ship, the persistence is half the baseline value. This is evidence of a strong shift in the behavior of the whales, which might have consequences on their capacity to rest and forage at short term.
@@ 7 Conclusion and perspectives
The authors would like to thank the CNRS and IRN Madef for funding A. Delporte, and MIAI - ANR-19-P3IA-0003 for the funding of A Samson. We also thank the project MATH-AmSud 23-MATH-12.